OpenStudy (anonymous):

how to solve sin^2phetacospheta=2sin^3pheta

OpenStudy (anonymous):

is this pheta or theta???

OpenStudy (ajprincess):

is ur question this? \(\sin^2\theta\cos\theta=2\sin^3\theta\) @amina.kalla

Directrix (directrix):

Let pheta = x. Then, assuming the equation is as written above by @ajprincess , we get: (sin² x) (cos x) = 2 sin³ x (sin² x) (cos x) - 2 sin³ x = 0 Factor out (sin² x) to get: (sin² x) [ cos x - 2 sin x ] = 0 Use the Zero Product Property to write: (sin² x) = 0 or [ cos x - 2 sin x ] = 0 @amina.kalla Your task is to solve the two equations directly above.

OpenStudy (anonymous):

thnks

Directrix (directrix):

Did you get it solved? @amina.kalla

OpenStudy (anonymous):

can u help me with this quest cos(theta)=sin(theta/2)

Directrix (directrix):

I will need to look up the half angle formula for sine.

OpenStudy (anonymous):

cos(3theta)

OpenStudy (anonymous):

=sin(theta/2)

Directrix (directrix):

I thought this is the question: cos(theta)=sin(theta/2)

OpenStudy (anonymous):

the quest is cos (3theta)

Directrix (directrix):

@amina.kalla I don't know how to solve that trig equation now.

OpenStudy (ajprincess):

so the question is \(\cos3\theta=\sin(\theta/2)\). right?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

\(\sin(\theta/2)=\cos(\pi-(\theta/2))\) so \(\cos3\theta=\cos(\pi-(\theta/2))\) \(3\theta=\pi-(\theta/2)\) nw can u find \(\theta\)?

OpenStudy (anonymous):

ok thnks

OpenStudy (ajprincess):

welcome:)